# Binary VPN Traffic Detection Using Wavelet Features and Machine Learning

### v1

@@ Abstract
Encrypted traffic classification faces growing challenges as encryption renders traditional deep packet inspection ineffective. This study addresses binary VPN detection, distinguishing VPN-encrypted from non-VPN traffic using wavelet transform-based features across multiple machine learning models. We analyze the impact of wavelet decomposition levels and dataset filtering on classification performance. Our results demonstrate that Random Forest (RF) achieves superior performance with an F1-score of 99%, maintaining robust accuracy even after significant dataset filtering. Neural Networks (NN) show comparable effectiveness with an F1-score of 98% when trained on wavelet level 12, while Support Vector Machines (SVM) exhibit notable sensitivity to dataset reduction, with F1-scores dropping from 90% to 85% after filtering. Comparing wavelet decomposition at levels 5 and 12, we observe improved classification performance at level 12, particularly for variable traffic types, though the marginal gains may not justify the additional computational overhead. These findings establish RF as the most reliable model for VPN traffic classification while highlighting key performance tradeoffs in feature extraction and preprocessing.
@@ 1 Introduction
In this work, we focus on binary VPN detection and re-evaluate the effectiveness of wavelet transform-based features for distinguishing VPN from non-VPN traffic. We assess the performance of multiple ML models, including Random Forest (RF), Neural Networks (NN), and Support Vector Machines (SVM), and analyze the impact of dataset filtering, a common preprocessing step that significantly affects classification performance and data distribution. Unlike prior work that primarily targets application classification, our analysis provides insights into the robustness of different ML models under specific data processing variations: wavelet decomposition depth (levels 5 and 12) and flow-length filtering.
@@
[…]
2 Wavelet Transform
Wavelet transform is a powerful mathematical tool for analyzing non-stationary signals by decomposing them into components of varying frequency and localization in time. Unlike Fourier transforms, which provide only frequency-domain representation, wavelet transforms preserve both time and frequency information, making them particularly suited for analyzing complex and transient signals such as network traffic [Ahad2016]. This section outlines the fundamentals of wavelet features, their general applications, and their specific use in the analysis of IP flows for VPN detection.
2.1 Wavelet Transform Fundamentals
2.1.1 Types of Wavelet Transforms
Several variations of wavelet transforms exist, each optimized for specific applications:
  • Continuous Wavelet Transform (CWT): Provides a detailed time-frequency representation of signals by continuously scaling and shifting the wavelet function, making it ideal for analyzing non-stationary signals with high precision.
  • Discrete Wavelet Transform (DWT): Operates on discrete scales and positions, enabling efficient computation and perfect reconstruction of signals. It is particularly suited for tasks such as signal compression and noise reduction.
  • Wavelet Packet Transform (PWT): Extends DWT by further decomposing both approximation and detail coefficients, offering a finer resolution for signals with intricate frequency components [RuizdelaHermosaGonzlezCarrato2014].
Among these, DWT is preferred for its balance between analytical capability and computational efficiency.
2.1.2 Discrete Wavelet Transform Fundamentals
DWT decomposes a signal into two sets of coefficients: approximation coefficients, which capture the low-frequency components, and detail coefficients, which represent the high-frequency components of the signal [91217].
  • Approximation coefficients ($A_{j}$) are computed as follows:
    $A_{j}[k]=\sum_{n}x[n]\cdot g(2k-n),$ (1)
    where $x[n]$ is the input signal, $g[n]$ represents the low-pass filter coefficients, and $2k$ is the downsampling factor.
  • Detail coefficients ($D_{j}$) are computed as:
    $D_{j}[k]=\sum_{n}x[n]\cdot h(2k-n),$ (2)
    where $h[n]$ represents the high-pass filter coefficients.
Both approximation and detail coefficients are used to compute the wavelet-based features. To simplify the notation in feature calculations, we introduce a general notation $C_{j}[k]$, where $C_{j}[k]$ represents either $A_{J}[k]$ (approximation coefficient at the final level) or $D_{j}[k]$ (detail coefficients at intermediate levels).
2.1.3 Wavelet Transform Metrics
To analyze the coefficients $C_{j}[k]$ obtained from DWT, we compute several key metrics that characterize different aspects of the signal.
  1. 1.
    $E_{j}=\frac{\sum_{k}C_{j}[k]^{2}}{\sum_{j}\sum_{k}C_{j}[k]^{2}}\times 100\%.$ (3)
    Here, $\sum_{k}C_{j}[k]^{2}$ represents the energy of the coefficients at level $j$, while $\sum_{j}\sum_{k}C_{j}[k]^{2}$ represents the total energy across all levels [Amin2015]. Since this metric reflects the relative energy distribution, it satisfies the constraint $\sum_{j}E_{j}=100\%$.
  2. 2.
    Absolute Mean of Coefficients: The absolute mean ($M_{j}$) at level $j$ quantifies the average magnitude of wavelet coefficients [753747]:
    $M_{j}=\frac{1}{N_{j}}\sum_{k}|C_{j}[k]|.$ (4)
    Here, $N_{j}$ represents the number of coefficients at level $j$, and $|C_{j}[k]|$ is the absolute value of each coefficient.
  3. 3.
    Standard Deviation of Coefficients: The standard deviation ($\sigma_{j}$) at level $j$ measures the spread of coefficients [7566534]:
    $\sigma_{j}=\sqrt{\frac{1}{N_{j}}\sum_{k}\left(C_{j}[k]-\bar{C}_{j}\right)^{2}}.$ (5)
    where the mean $\bar{C}_{j}$ is given by:
    $\bar{C}_{j}=\frac{1}{N_{j}}\sum_{k}C_{j}[k].$ (6)
  4. 4.
    Shannon Entropy: The Shannon entropy ($H_{j}$) at level $j$ quantifies the information content of the wavelet coefficients [Uyar2008]:
    $H_{j}=-\sum_{k}P_{j}[k]\log_{2}P_{j}[k].$ (7)
    The probability distribution $P_{j}[k]$ is defined as:
    $P_{j}[k]=\frac{|C_{j}[k]|}{\sum_{k}|C_{j}[k]|},$ (8)
    with the constraint:
    $\sum_{k}P_{j}[k]=1.$ (9)
2.2 Wavelet Features in IP Flows
IP flows consist of sequences of packets sharing common attributes (e.g., source/destination IP addresses and ports, and protocols) within a specific time window. Wavelet-based analysis is applied to numerical sequences derived from flow attributes such as packet size.
2.2.1 Decomposition Process
The DWT decomposes these sequences into detail coefficients ($D_{j}$) and approximation coefficients ($A_{j}$) over multiple decomposition levels ($J$). The optimal number of decomposition levels is determined as:
$J=\lfloor\log_{2}(n)\rfloor,$ (10)
where $n$ is the length of the sequence.
2.2.2 Feature Vector Construction
For each decomposition level, wavelet features are computed to create comprehensive feature vectors for classification. These include:
  • Energy Features: Raw energy values ($E_{j}$) at each decomposition level.
  • Statistical Features:
    • Absolute mean ($M_{j}$).
    • Standard deviation ($\sigma_{j}$).
    • Shannon entropy ($H_{j}$).
@@ 2.2.2 Feature Vector Construction
Complete Flows Act. Timeout 41 Act. Timeout 4.96
@@
3 Methodology
This section describes our approach to VPN traffic classification. We first present the dataset and our flow metering process, followed by feature extraction methodology for the wavelet-based features. We then detail the machine learning models employed and the evaluation metrics used to assess their performance.
3.1 Dataset Description
@@
3.2 Flow Processing and Analysis
We utilized NFStream [Aouini2022], a flexible network data analysis framework, to process the dataset’s PCAP files. While VNAT provides both traffic traces in PCAP format and pre-extracted statistical summaries in HDF (h5) format, we chose NFStream for its ability to seamlessly integrate plugins, particularly for calculating the Wavelet Features discussed in Section 2.
To ensure consistency with the methodology described in [10044382], we extracted flow records under two conditions: first, without applying any flow timeout to capture complete flow records, and second, with an active timeout of 41 seconds, approximating the original study’s segmentation of flows into 40.96-second intervals. In the latter case, packets were assigned to 0.01-second time bins, and intervals containing fewer than 20 packets were discarded.
@@ 3.2 Flow Processing and Analysis
The reduction in flow counts is particularly pronounced in Command & Control (91.64%), File Transfer (95.52%), and VoIP (80.21%) traffic. This dramatic reduction highlights the effect of flow segmentation on dataset composition and raises potential concerns regarding class imbalance, as we will discuss in Section 5.
@@
3.3 Feature Extraction
For our classification task, we extract wavelet-based features to capture the temporal characteristics of packet size distributions in network flows. We apply DWT to transport-layer packet size sequences in both forward and backward directions. Following [10044382], we remove flows with fewer than 20 packets. Our choice of decomposition level J=5 is supported by the packet count distribution in our filtered dataset, as shown in Table II. With the 25th percentile at 56 packets, J=5 (scale of $322^{5}=$ 32 via Equation 10) ensures effective capture of packet variations across the majority of flows. We also evaluate J=12 to enable direct comparison with [10044382], though at this level (scale of $40962^{12}=$ 4096), the decomposition scale significantly exceeds typical flow lengths in their dataset. The highly skewed distribution (mean $=2,599$, median $=63$) indicates that while some flows contain up to 857,900 packets, such extensive flows are rare outliers. This distribution suggests that while J=12 might capture finer-grained variations in exceptionally long flows, it might provide diminishing returns for typical traffic patterns.
@@ 3.3 Feature Extraction
From each decomposition, we compute four key wavelet-based metrics using both the approximation coefficients Equation 1 and detail coefficients Equation 2: relative wavelet energy Equation 3, absolute mean Equation 4, standard deviation Equation 5, and Shannon entropy Equation 7. These metrics quantify the distribution of energy across frequency components, capturing structural properties of packet sequences.
@@
TABLE III: Performance Comparison of Machine Learning Models
@@
3.4 Classification Models
3.4.1 Random Forest
RF is implemented as an ensemble-based classifier, leveraging decision tree aggregation to improve generalization. It is well-suited for high-dimensional feature spaces and has been widely used in network traffic classification due to its robustness and interpretability. Default hyperparameters were used, as preliminary evaluation showed competitive performance across the dataset without requiring extensive optimization.
3.4.2 Neural Networks
3.4.3 Support Vector Machine
Figure 1: Accuracy Comparison of ML Models with and without Filtering.
3.5 Performance Metrics
Beyond classification metrics, we investigate the effect of filtering on model performance. Since filtering reduces dataset size by discarding flows with fewer than 20 packets, it may introduce biases or reduce model generalizability.The impact of filtering on different models will be discussed in Section 4.
@@ 3.5 Performance Metrics
Figure 3: Distribution of Misclassified Samples Across Different Traffic Categories for RF 12 levels
4 Results
Table III summarizes our experimental results. For visual comparison, we provide the accuracy scores separately as bar plots in Fig. 1. Models are denoted with suffixes ‘5’ or ‘12’ indicating the wavelet decomposition levels used for feature extraction, while the suffix ‘_filtered’ identifies models trained on datasets where, following the methodology in [10044382], flows with fewer than 20 packets were removed.
@@ 4 Results
From Table III, we observe that RF maintains consistently high F1-scores of 99% at both decomposition levels, with only a minor decrease to 98% after filtering. NNs exhibit sensitivity to decomposition levels, with NN12 achieving an F1-score of 98%, slightly decreasing to 96% after filtering, while NN5 remains stable at 93% across both filtered and unfiltered scenarios. SVMs show the most significant impact from filtering. The F1-score for SVM12 drops from 90% to 85%, while SVM5 declines from 88% to 83% after filtering, indicating that SVM remains the most affected model by dataset reduction.
@@
The accuracy patterns in Fig. 1 align with these observations, with RF maintaining stability at both decomposition levels, showing only minimal reductions from 99% to 98% after filtering. NN demonstrates moderate sensitivity, with NN12 accuracy decreasing from 98% to 96%, while NN5 remains unchanged at 93% post-filtering. SVM exhibits the most noticeable filtering impact, with SVM12 accuracy declining from 90% to 85% and SVM5 dropping from 88% to 83%, reinforcing that SVM is the most sensitive model to dataset reduction.
@@
Since RF achieved consistently the best performance across both 5 and 12-level decomposition configurations, we focus our detailed analysis on its confusion matrices and misclassifications.
@@
Fig. 2 illustrates the classification performance across different RF configurations. RF12 achieves 100% accuracy for non-VPN traffic and 97% for VPN traffic, with minimal misclassification rates of 0% and 3%, respectively. The filtered version maintains strong performance, with 99% accuracy for non-VPN and 98% for VPN traffic, showing a slight reduction in misclassification rates. RF5 exhibits similar performance, achieving 100% accuracy for non-VPN and 97% for VPN traffic. After filtering, RF5 shows a small drop in non-VPN classification accuracy to 98%, while VPN accuracy improved to 98%. These results confirm RF’s robustness, with only minimal variations in misclassification rates across different wavelet levels and filtering conditions.
@@
Fig. 3 reveals category-specific classification challenges. Chat traffic presents the highest misclassification counts, with RF12 showing 46 total misclassifications (17 non-VPN, 29 VPN) before filtering, reducing to 35 (20 non-VPN, 15 VPN) after filtering. File Transfer and Command & Control categories show marked improvement with filtering—RF12’s misclassifications in File Transfer drop from 8 to 3 cases. RF5 shows similar patterns but with generally higher misclassification counts across all categories, particularly visible in the Chat and Streaming categories, reinforcing the advantage of higher decomposition levels for accurate classification.
However, these reduced misclassification counts after filtering should be interpreted with caution. As shown in Table I, filtering dramatically reduces the available samples—up to 95.52% in File Transfer and 91.64% in Command & Control categories. Thus, the lower misclassification counts primarily reflect the reduced dataset size rather than actual performance improvement. This observation aligns with the accuracy and F1-score metrics in Table III, which show that filtering generally decreases model performance despite the apparent reduction in misclassifications.
5 Discussion
RF outperforms NN and SVM in VPN classification across all tested conditions. RF’s ensemble structure handles missing information in shorter flows effectively, while SVM degrades due to its sensitivity to feature sparsity and distribution shifts.
The superior performance of RF and NN with J=12 over J=5 highlights the benefits of finer-grained spectral features, particularly in Chat and Streaming traffic. However, RF12 shows diminishing returns over RF5, suggesting that beyond a certain threshold, deeper wavelet decomposition adds little value. Filtering disproportionately reduces File Transfer and Command & Control traffic, exacerbating class imbalance and increasing misclassification risks, especially in underrepresented categories.
Misclassification analysis shows Chat traffic has the highest error rate, likely due to overlapping encrypted and non-encrypted patterns, while VoIP remains perfectly classified by RF, indicating distinct traffic characteristics.
Beyond accuracy, these findings highlight the importance of structured feature spaces in encrypted traffic analysis. NN models’ sensitivity to dataset variations suggests benefits from pretraining or feature augmentation. Future work could explore adaptive wavelet decomposition, focusing on the most informative spectral bands to improve efficiency and scalability.
6 Comparison with VNAT Study

### v2

@@ Abstract
Encrypted traffic classification faces growing challenges as encryption renders traditional deep packet inspection ineffective. This study addresses binary VPN detection, distinguishing VPN-encrypted from non-VPN traffic using wavelet transform-based features across multiple machine learning models. Unlike previous studies focused on application-level classification within encrypted traffic, we specifically evaluate the fundamental task of VPN identification regardless of application type. We analyze the impact of wavelet decomposition levels and dataset filtering on classification performance across significantly imbalanced data, where filtering reduces some traffic categories by up to 95%. Our results demonstrate that Random Forest (RF) achieves superior performance with an F1-score of 99%, maintaining robust accuracy even after significant dataset filtering. Neural Networks (NN) show comparable effectiveness with an F1-score of 98% when trained on wavelet level 12, while Support Vector Machines (SVM) exhibit notable sensitivity to dataset reduction, with F1-scores dropping from 90% to 85% after filtering. Comparing wavelet decomposition at levels 5 and 12, we observe improved classification performance at level 12, particularly for variable traffic types, though the marginal gains may not justify the additional computational overhead. These findings establish RF as the most reliable model for VPN traffic classification while highlighting key performance tradeoffs in feature extraction and preprocessing.
@@ I Introduction
In this work, we focus on binary VPN detection and re-evaluate the effectiveness of wavelet transform-based features for distinguishing VPN from non-VPN traffic. We assess the performance of multiple ML models, including Random Forest (RF), Neural Networks (NN), and Support Vector Machines (SVM), and analyze the impact of dataset filtering, a common preprocessing step that significantly affects classification performance and data distribution. Our analysis deliberately includes both filtered and unfiltered datasets to evaluate the robustness of models across varying conditions, even with the highly imbalanced class distributions present in real-world traffic. Unlike prior work that primarily targets application classification, our analysis provides insights into the robustness of different ML models under specific data processing variations: wavelet decomposition depth (levels 5 and 12) and flow-length filtering.
@@
The remainder of this paper is structured as follows: Section II describes our methodology, including dataset preprocessing and model selection. Section III presents and analyzes our results. Finally, Section IV concludes this paper.
@@
Complete Flows Act. Timeout 41 Act. Timeout 40.96
@@
II Methodology
II-A Dataset Description
@@
II-B Flow Processing and Analysis
We utilized NFStream [Aouini2022], a flexible network data analysis framework, to process the dataset’s PCAP files. While VNAT provides both traffic traces in PCAP format and pre-extracted statistical summaries in HDF (h5) format, we chose NFStream for its ability to seamlessly integrate plugins, particularly for calculating the Wavelet Features discussed in Section II-C.
To ensure consistency with the methodology described in [10044382], we extracted flow records under two conditions: first, without applying any flow timeout to capture complete flow records, and second, with an active timeout of 41 seconds, approximating the original study’s segmentation of flows into 40.96-second intervals. In the latter case, packets were assigned to 0.01-second time bins, and intervals containing fewer than 20 packets were discarded. This dual approach enables us to evaluate the impact of flow segmentation and filtering on classification performance.
@@ II-B Flow Processing and Analysis
The reduction in flow counts is particularly pronounced in Command & Control (91.64%), File Transfer (95.52%), and VoIP (80.21%) traffic. This dramatic reduction highlights the effect of flow segmentation on dataset composition and raises potential concerns regarding class imbalance.
@@
This filtering has a non-uniform effect on VPN and non-VPN traffic distributions, altering the class balance within categories and creating a more challenging classification scenario that better reflects real-world deployment conditions.
II-C Wavelet Features
Wavelet transform is a time-frequency analysis method that decomposes signals into components at different scales, making it suitable for analyzing non-stationary traffic patterns. Unlike Fourier transforms, wavelets preserve both time and frequency information, capturing traffic bursts and periodic behaviors effectively [Ahad2016].
The Discrete Wavelet Transform (DWT) decomposes a signal into approximation coefficients (low-frequency components) and detail coefficients (high-frequency components) at multiple levels. For a network flow, the signal is the sequence of packet sizes over time [91217]. The approximation coefficients are computed as:
$A_{j}[k]=\sum_{n}x[n]\cdot g(2k-n)$ (1)
where $x[n]$ is the input signal, $g[n]$ represents the low-pass filter coefficients, and $2k$ is the downsampling factor. The detail coefficients are computed as:
$D_{j}[k]=\sum_{n}x[n]\cdot h(2k-n)$ (2)
where $h[n]$ represents the high-pass filter coefficients.
From these coefficients, we extract four key metrics that characterize different aspects of network traffic:
  1. 1.
    $E_{j}=\frac{\sum_{k}C_{j}[k]^{2}}{\sum_{j}\sum_{k}C_{j}[k]^{2}}\times 100\%$ (3)
  2. 2.
    Absolute Mean of Coefficients ($M_{j}$) quantifies the average magnitude at level $j$  [753747]:
    $M_{j}=\frac{1}{N_{j}}\sum_{k}|C_{j}[k]|$ (4)
  3. 3.
    Standard Deviation of Coefficients ($\sigma_{j}$) measures coefficient spread [7566534]:
    $\sigma_{j}=\sqrt{\frac{1}{N_{j}}\sum_{k}\left(C_{j}[k]-\bar{C}_{j}\right)^{2}}$ (5)
  4. 4.
    Shannon Entropy ($H_{j}$) quantifies the information content [Shannon1948]:
    $H_{j}=-\sum_{k}P_{j}[k]\log_{2}P_{j}[k]$ (6)
    where $P_{j}[k]=\frac{|C_{j}[k]|}{\sum_{k}|C_{j}[k]|}$.
The optimal number of decomposition levels is determined as $J=\lfloor\log_{2}(n)\rfloor$, where $n$ is the sequence length. This provides a mathematical basis for our choice of decomposition levels in feature extraction.
II-D Feature Extraction
For our classification task, we extract wavelet-based features to capture the temporal characteristics of packet size distributions in network flows. We apply DWT to transport-layer packet size sequences in both forward and backward directions. Following [10044382], we remove flows with fewer than 20 packets. Our choice of decomposition level J=5 is supported by the packet count distribution in our filtered dataset, as shown in Table II. With the 25th percentile at 56 packets, J=5 (scale of $322^{5}=$ 32 according to the optimal decomposition level formula $J=\lfloor\log_{2}(n)\rfloor$) ensures effective capture of packet variations across the majority of flows. We also evaluate J=12 to enable direct comparison with [10044382], though at this level (scale of $40962^{12}=$ 4096), the decomposition scale significantly exceeds typical flow lengths in their dataset. The highly skewed distribution (mean $=2,599$, median $=63$) indicates that while some flows contain up to 857,900 packets, such extensive flows are rare outliers. This distribution suggests that while J=12 might capture finer-grained variations in exceptionally long flows, it might provide diminishing returns for typical traffic patterns.
@@ II-D Feature Extraction
From each decomposition, we compute four key wavelet-based metrics using both the approximation coefficients and detail coefficients: relative wavelet energy, absolute mean, standard deviation, and Shannon entropy. These metrics quantify the distribution of energy across frequency components, capturing structural properties of packet sequences.
@@
II-E Classification Models
II-E1 Random Forest
RF is implemented as an ensemble-based classifier, leveraging decision tree aggregation to improve generalization. Default hyperparameters were used, as preliminary evaluation showed competitive performance across the dataset without requiring extensive optimization.
II-E2 Neural Networks
II-E3 Support Vector Machine
II-F Performance Metrics
Beyond classification metrics, we investigate the effect of filtering on model performance. Since filtering reduces dataset size by discarding flows with fewer than 20 packets, it may introduce biases or reduce model generalizability. We deliberately evaluate performance on both filtered and unfiltered datasets to determine how robust each model is to variations in flow length, an important consideration for real-world deployment where flows of all lengths will be encountered. Section III discusses the impact of filtering on different models.
Analyzing both filtered and unfiltered datasets, despite the predominance of very short flows (1-2 packets) in the latter, serves two important purposes: first, it establishes a performance baseline for comparison across preprocessing techniques; second, it evaluates the models’ capability to extract meaningful signal from minimal data, which is essential in scenarios where early traffic detection is required before sufficient packets have been observed.
III Results
Table III summarizes our experimental results across all model configurations. For visual comparison, we provide the accuracy scores as bar plots in Fig. 1. Models are denoted with suffixes ‘5’ or ‘12’ indicating the wavelet decomposition levels used for feature extraction, while the suffix ‘_filtered’ identifies models trained on datasets where flows with fewer than 20 packets were removed.
TABLE III: Performance Comparison of ML Models
@@ III Results
Figure 1: Accuracy Comparison with and without Filtering
@@
Figure 3: Distribution of Misclassified Samples Across Different Traffic Categories for RF 5 and 12 levels
@@
From Table III, we observe that RF maintains consistently high F1-scores of 99% at both decomposition levels, with only a minor decrease to 98% after filtering. This remarkable stability across both preprocessing conditions demonstrates RF’s robust feature extraction capability even with minimal data. NNs exhibit sensitivity to decomposition levels, with NN12 achieving an F1-score of 98%, slightly decreasing to 96% after filtering, while NN5 remains stable at 93% across both filtered and unfiltered scenarios. This 5% performance gap between NN12 and NN5 suggests that NNs benefit from the additional spectral information in higher decomposition levels. SVMs show the most significant impact from filtering, with F1-scores for SVM12 dropping from 90% to 85% (5% decrease), while SVM5 declines from 88% to 83% (also 5% decrease), indicating consistent sensitivity to dataset reduction regardless of decomposition level. This sensitivity likely stems from SVM’s reliance on support vectors at the class boundary, which become more sparse after filtering.
@@
The accuracy patterns in Fig. 1 confirm these observations. RF maintains high stability at both decomposition levels, showing only minimal reductions from 99% to 98% after filtering. NN demonstrates moderate sensitivity, with NN12 accuracy decreasing from 98% to 96%, while NN5 remains unchanged at 93% post-filtering. SVM exhibits the most pronounced filtering impact, with SVM12 accuracy declining from 90% to 85% and SVM5 dropping from 88% to 83%.
@@
Since RF achieved consistently the best performance across both decomposition configurations, we focus our detailed analysis on its confusion matrices and misclassification patterns.
@@
Fig. 2 illustrates the classification performance across different RF configurations. RF12 achieves 100% accuracy for non-VPN traffic and 97% for VPN traffic, with minimal misclassification rates of 0% and 3%, respectively. The filtered version maintains strong performance, with 99% accuracy for non-VPN and 98% for VPN traffic. RF5 exhibits similar performance, achieving 100% accuracy for non-VPN and 97% for VPN traffic. After filtering, RF5 shows a small drop in non-VPN classification accuracy to 98%, while VPN accuracy improved to 98%. These results confirm RF’s robustness, with only minimal variations in misclassification rates across different wavelet levels and filtering conditions.
@@
Fig. 3 reveals category-specific classification challenges. Chat traffic presents the highest error counts, with RF12 showing 46 total misclassifications (17 non-VPN, 29 VPN) before filtering, reducing to 35 (20 non-VPN, 15 VPN) after filtering. When viewed relative to category size (1301 Chat samples), this represents approximately 3.5% of Chat samples misclassified. The higher error rate likely stems from Chat’s variable traffic patterns and similarities between encrypted and non-encrypted protocols. File Transfer and Command & Control categories show marked improvement with filtering—RF12’s misclassifications in File Transfer drop from 8 to 3 cases, suggesting that longer flows contain more distinctive VPN signatures. RF5 shows similar patterns but with generally higher misclassification counts across all categories, particularly visible in the Chat and Streaming categories.
Notably, as shown in Fig. 3, VoIP traffic achieves zero misclassifications in the RF12 model under both filtered and unfiltered conditions, indicating highly distinctive traffic patterns between VPN and non-VPN variants despite the 80% dataset reduction from filtering. This exceptional performance suggests that VoIP applications maintain unique temporal signatures despite encryption, likely due to their real-time streaming requirements.
When interpreting the reduced misclassification counts after filtering, it’s important to consider the dramatic reduction in available samples—up to 95.52% in File Transfer and 91.64% in Command & Control categories, as shown in Table I. Thus, while absolute misclassification counts decrease after filtering, the error rates as a percentage of category size actually increase slightly in several categories (File Transfer error rate increases from 0.05% to 0.37% after filtering). This observation aligns with the overall F1-score patterns in Table III, confirming that filtering generally causes a slight performance degradation.
The performance comparison between J=12 and J=5 reveals a consistent advantage for deeper wavelet decomposition, with improvements ranging from 1% (RF12 vs. RF5) to 5% (NN12 vs. NN5) in F1-scores. However, these relatively modest gains suggest diminishing returns that may not justify the increased computational overhead in all deployment scenarios, particularly in resource-constrained environments where model efficiency is prioritized alongside accuracy.
IV Conclusion
This study evaluated binary VPN traffic classification using wavelet-based features and machine learning models, focusing on the impact of wavelet decomposition levels and dataset filtering. Our results show that RF consistently achieves the highest classification accuracy, maintaining robustness across different preprocessing conditions, while SVM exhibits the greatest sensitivity to dataset reductions. We found that deeper wavelet decompositions improve classification performance, particularly for variable traffic types, though the marginal gains may not justify the additional computational overhead. Future research could explore adaptive wavelet decomposition techniques that dynamically adjust to flow characteristics and investigate ensemble approaches combining multiple decomposition levels for enhanced robustness across diverse traffic conditions.
This work utilizes the VNAT dataset [10044382] as the primary data source. All processed datasets derived from these traces—including extracted network flows, filtered datasets, and wavelet-based feature representations—along with analysis scripts, trained models, and performance metrics necessary to reproduce our results are available at [github-repo], ensuring full reproducibility of the machine learning experiments.
Acknowledgement
Supported by the János Bolyai Research Scholarship of the Hungarian Academy of Sciences. This work was also supported by project TKP2021-NVA-02 (financed under the TKP2021-NVA scheme), implemented with support provided by the Ministry of Culture and Innovation of Hungary from the National Research, Development and Innovation Fund.